The lengths of the sides of a triangle are 3, 4, and 6. If the length of the shortest side of a similar triangle is 5, find the length of its longest side.
Answer by TimothyLamb(4379)

for similar triangles, corresponding sides have the same ratio:
---
5/3 = x/6
x = 6*5/3
x = 10
---
answer:
the longest side of the new triangle is 10
---
Solve and graph linear equations:
https://sooeet.com/math/linear-equation-solver.php
---
Solve quadratic equations, quadratic formula:
https://sooeet.com/math/quadratic-formula-solver.php
---
Convert fractions, decimals, and percents:
https://sooeet.com/math/fraction-decimal-percent.php
---
Calculate and graph the linear regression of any data set:
https://sooeet.com/math/linear-regression.php